Isis can paint her garage in 1 hour and 15 minutes. If Jerry helps her, they can paint the garage in 45 minutes. How long would it take Jerry to paint the garage by himself?
Answer by nerdybill(7384)

Isis can paint her garage in 1 hour and 15 minutes. If Jerry helps her, they can paint the garage in 45 minutes. How long would it take Jerry to paint the garage by himself?
1 hour and 15 minutes = 60 + 15 = 75 minutes
.
Let x = time (minutes) it takes Jerry to paint alone
then
45(1/75 + 1/x) = 1
multiplying both sides by 75x:
45(x + 75) = 75x
45x + 3375 = 75x
3375 = 30x
112.5 minutes = x
or
1 hour and 52.5 minutes = x